OpenStudy (anonymous):

What is the sum of 1/2 + 1/2^2 + 1/2^3 +...+ 1/2^2012? Help me please?

OpenStudy (zarkon):

let S= 1/2 + 1/2^2 + 1/2^3 +...+ 1/2^2012 then 2S= 1+ 1/2 + 1/2^2 +...+ 1/2^2011 subtract the two equations